OpenStudy (anonymous):

CAN ANY BODY ANSWER IT how many ways can 6 coins be chosen from 20 one rupee coins,10 fifty paise coins,7 twenty paise coins...

OpenStudy (anonymous):

In total, you have 20 + 10 + 7 = 37 coins. If the type of coin doesn't matter, then you have \[_{37}P_6=\frac{37!}{(37-6)!}=\frac{37!}{31!}=37\cdot36\cdot35\cdot34\cdot33\cdot32=1,673,844,480.\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

but the answer is 28

OpenStudy (kropot72):

@msingh What are the full details of the question?

OpenStudy (anonymous):

this is the the only question

OpenStudy (anonymous):

how many ways can 6 coins be chosen from 20 one rupee coins,10 fifty paise coins,7 twenty paise coins.

OpenStudy (anonymous):

Do they have to add up to a certain amount?

OpenStudy (kropot72):

Given that the answer is 28 ways, it must be necessary to find the constraints on the selection of coins that result in the given answer.

OpenStudy (anonymous):

k so how it can be solved

OpenStudy (anonymous):

Hey @SithsAndGiggles ... I think the coins of different denominations are of course different...and that's why the question says "coins from 20 one rupee coins,10 fifty paise coins,7 twenty paise coins..." Had it not been the case, the question would have told 37 coins are there... Between, @msingh , I would give the explanation tomorrow... already late... sorry... No worries...

OpenStudy (anonymous):

but i need it today

OpenStudy (phi):

they must mean that coins of the same denomination are indistinguishable There must be a nice formula, but brute force will answer this question. If we label the 3 types A B C and list the possible combinations: 6 0 0 5 0 1 5 1 0 4 0 2 4 1 1 4 2 0 and so on... you will get 28 combinations